The velocity of a blood corpuscle in a vessel depends on how far the corpuscle is from the center of the vessel. Let R be the constant radius of the vessel Vm, the constant maximum velocity of the corpuscle; r, the distance from a center to a particular blood corpuscle (variable); and Vr, the velocity of that corpuscle. The velocity Vr is related to the distance r
according to Vr=Vm(1-r2/R2). find f when Vr = 1/4 Vm

Answer by Alan3354(69443)

The velocity of a blood corpuscle in a vessel depends on how far the corpuscle is from the center of the vessel. Let R be the constant radius of the vessel Vm, the constant maximum velocity of the corpuscle; r, the distance from a center to a particular blood corpuscle (variable); and Vr, the velocity of that corpuscle. The velocity Vr is related to the distance r
according to Vr=Vm(1-r2/R2). find f when Vr = 1/4 Vm
-----------------------
What is f?